# Reflection on Data Storytelling Tools in the Generative AI Era from the Human-AI Collaboration Perspective

### v1

@@ 4.1. What pattern remains popular?
The most popular patterns with both humans’ and AI’s participation in earlier tools are human-creator + AI-assistant  [^2] The plus sign indicates the two roles collaborate with each other in a specific stage. and AI-creator + human-optimizer since they allow humans to control the content while lowering the barrier of creation through automation (Li et al., 2024b). In the latest tools, their frequency only has limited fluctuation, suggesting that they are still welcomed. We noticed two noteworthy advancements brought by generative AI models when applying the patterns.
First, in the communication stage, we identify an innovative usage of AI-assistant, which may lead to future research opportunities. The usage of AI-assistant in the communication stage was to support individual users in live presentations, such as retrieving and presenting charts based on sketch (Lee et al., 2013) and supporting interaction with charts with gesture recognition (Hall et al., 2022). Building on LLMs’ ability to understand and generate texts, Han and Isaacs (Han and Isaacs, 2024) proposed an AI-assistant to record the communication process between human-creators around charts and create an interactive data document with texts, visualizations, and annotations from the conversation. The work has more implications beyond identifying a new function of AI-assistants powered by generative AI models. It demonstrates a new collaborative data story creation mode, where humans freely express their insights around data in conversations and AI documents and organizes humans’ ideas into a more structured data story. It is interesting to examine the benefit of this collaboration mode and explore other potentials. For example, will it be possible to ask AI to interview humans and write data stories or generate engaging data videos like talk shows based on humans’ conversations around data? Furthermore, from the human-AI collaboration perspective, this AI-assistant has a notable difference with others. It serves as a third-party collaborator to document multiple humans’ communication, while not collaborating with individual humans and contributing to their creation or presentation (e.g., AI-assistants in VizFlow (Sultanum et al., 2021) and SketchStory (Lee et al., 2013)). We should further investigate the potentials of human-AI collaboration to support multiple humans’ work in data storytelling (Chevalier et al., 2018).
@@
Second, the application of large-scale generative AI can improve the collaboration experience from several perspectives. The first perspective is to increase AI collaborators’ output space significantly. In our corpus, we notice that Epigrahics (Zhou et al., 2024a), AnalogyMate (Chen et al., 2024), and ChartSpark (Xiao et al., 2023) follow a similar design in the implementation stage. Their AI-assistants apply generative models to create visual components and provide design inspirations to support human-creators for authoring static or animated infographics. The similar design exists in earlier tools, such as Vistylist (Shi et al., 2022), Chartreuse (Cui et al., 2021) and DataQuilt (Zhang et al., 2020). However, they can only support re-using existing visual elements, greatly limiting the potential space of creation. The AI collaborators powered by retrieval-based algorithms also suffer from the issue , such as AI-creators in Retrieve-then-Adapt (Qian et al., 2020) and Text-to-Viz (Cui et al., 2019). Compared to these earlier work, the latest tools powered by generative AI models are more flexible to create new visual elements beyond existing examples. Additionally, the interaction between collaborators can be more natural and closer to humans’ habits. Besides direct text instructions (Xiao et al., 2023; Zhao et al., 2024), OutlineSpark (Wang et al., 2024a) allows humans to input a structured outline which guides AI-creator to select cells from notebooks and generate structured slides. In Data Playwright (Shen et al., 2024b), users write annotated narrations to specify expected animation effect and timing along with narrations. In both tools, users’ intent are conveniently expressed to enhance their agency, and their efforts are reduced through more AI automation. The last advantage is output quality improvement. For example, LLMs are applied to enhance template-based text generation (Lin et al., 2023; Kim et al., 2024).
4.2. What pattern has newly appeared?
@@
We are happy to see that there are new collaboration pattern appearing in the latest tools: the usage of human-reviewers in the analysis and the planning stages and AI-creators in the communication stage.
@@
The collaboration pattern of AI-creator + human-reviewer has been discussed as a potential strategy to enhance AI automation and human agency (Li et al., 2024b). Compared to human-optimizers, human-reviewers can control the data story with feedback on AI-created content and thus save the effort to revise the content manually. In our surveyed tools, Socrates (Wu et al., 2023) allows human-reviewer to provide feedback to the direction of data analysis (e.g., what attributes in datasets to be included in the data story) and the narrative structure (e.g., highlight a data fact using the strategy of contrasting it with another one). Specifically, the AI-creator, powered by a heuristic-based algorithm, will generate questions and ask for human-reviewer to rate, compare, or select options. Then, the AI-creator can leverage the results to create a data story that aligns with the intent of human-reviewer. Though they do not explicitly compare the efforts spent with their tool and another with human-optimizer, i.e., Calliope (Shi et al., 2020), their results show that the created stories with Socrates match the intent better, which also implies that humans might save efforts for further adjustment. The study results preliminarily verify the advantages of human-reviewer to achieve more AI automation and more human agency than human-optimizer, when collaborating with an AI-creator. In addition to the authors’ suggestion of exploring free-form feedback, future research can further explore the design and application of human-reviewers, such as understanding the pros and cons explicitly and providing feedback to visual channels in data stories.
@@
Another new pattern is the application of AI-creators for communicating data stories automatically. As revealed in a previous interview study (Li et al., 2023a), data workers often prefer not to ask AI to perform the creator role in the communication since they may diminish the underlying value of human-human communication, such as building trust. Therefore, we are particularly interested in the design of these AI-creators in the latest tools. Among the four AI-creators, Sportify (Lee et al., 2024) and AiCommentator (Andrews et al., 2024) leverage AI-creators to communicate automatically identified data insights in sports games to humans. Ying et al. (Ying et al., 2024) and Shi et al. (Shi et al., 2024) support chart and dataset understanding with AI-creators who communicate insights to humans with animations and narrations. Comparing these cases with the interview study, we can notice that the AI-creator communicates AI-created stories in these tools rather than presenting human-created stories as their proxies. It implies whether AI-creators can be accepted by humans in the communication stage may depend on AI’s roles in the previous stages. If the story is fully created by the AI collaborator in previous stages and reflect little human opinions, the role of AI-creators is more acceptable. On the other hand, when the story serves for the purpose of human-human communication and carries intensive information from human storytellers, the usage of AI-creators as a human proxy might raise concerns. Beyond its implication to AI-creators, the observation also has reveals that the application of different roles across stages can be dependent on each other. It can be interesting to further understand the factors that lead to the dependency and explore whether the dependency of roles exists in other domains.
4.3. What pattern is getting more popular?
@@
Across all four stages, we notice the growing interest of the AI-creator-only pattern, where various fully automatic tools across multiple stages emerge (Ying et al., 2024; Lee et al., 2024; Shen et al., 2024a; Andrews et al., 2024). We consider that the observation is led by the capability of large-scale generative models, such as the power of analyzing data (Inala et al., 2024) and generating corresponding text explanation (Dibia, 2023). One might question if the trend contradicts with the argument that human-AI collaboration in data storytelling would gain more interests (Li et al., 2024b). Furthermore, does the advancement of AI capability mean that we should try to automate everything with AI in data storytelling and thus the research on human-AI collaboration is no longer needed?
@@
Inspired by the discussion where collaboration patterns can be affected by tool usage scenarios (Li et al., 2024b), we dive into the scenarios of these fully automatic tools to figure out why they are designed to be AI-creators-only. In the usage scenarios of Sportify (Lee et al., 2024) and AiCommentator (Andrews et al., 2024), sports fans issue targeted queries about matches for desired information, such as tactics or player performances. Then, AI collaborators create tailored and casual data stories based on the queries and communicate them to humans automatically. Compared to other scenarios, this one emphasizes the need of being customized and real-time while tolerating that the quality might not be perfect due to the casual nature of the scenario. When telling data stories to the public (Segel and Heer, 2010), the goal is often set by the authors and thus the stories are not tailored for a reader. In the scenario of Sportify and AiCommentator, data stories are told based on an individual user’s intent, implying that they have to be customized. Furthermore, making serious data stories for the public requires considerable time to polish them (Chevalier et al., 2018), while real-time processing might be prioritized in casual and low-stake scenarios (Lee et al., 2024; Andrews et al., 2024).
@@
To fulfill the requirements of customization and high responsiveness, the creation and communication of data stories can hardly be intervened by human-creators or human-assistants, who are likely to take longer time and higher financial cost to finish tasks (Li et al., 2023a). The traditional AI algorithms may also have limited flexibility for those highly customized data stories. The emergence of the latest large-scale AI models can fill the gap due to their ability of consuming free-form input to describe audiences’ goals and generating and communicating data stories (Section 4.1). They can serve as AI-creators to automate the whole workflow. Furthermore, since the scenarios are casual and low-stake, careful examination and quality improvement by human-optimizers or human-reviewers might not be needed.
@@
The scenario of Live Charts (Ying et al., 2024) and the work by Shi et al. (Shi et al., 2024) impose similar requirements to storytellers. The authors propose to animate static charts and support them with narration to facilitate understanding. The generated data stories should help user to interpret promptly but their designs do not need to be polished. Considering the high cost of humans further, a fully automatic pipeline is more suitable. Another similar audience-oriented scenario is to enhance text-only data story understanding by generating visuals (Li et al., 2024b), where automatic tools (Hullman et al., 2013; Gao et al., 2014; Shao et al., 2025) are designed.
@@
After examining these cases, we can attempt to answer the question about the necessity and potential future research direction of human-AI collaboration. First, led by the outstanding capability of large-scale generative AI models, it is a natural and irresistible trend that the application of AI-creators will be wider. To cope with the trend, we should actively experiment with these models (Li et al., 2024b) and learn how to apply them smartly, such as effective prompt design (Shen et al., 2024a). Furthermore, we should realize that the application of AI-creators is not only related to their performance but the scenario as well. The scenarios of fully automatic tools share common features, such as low-stake scenarios that require real-time generation of personalized stories. Additionally, there can be considerations regarding human-human communication (Section 4.2). In other scenarios, fully automatic tools might not be the suitable choice. For example, in contrast to the aforementioned tools (Lee et al., 2024; Andrews et al., 2024), SNIL (Cheng et al., 2024) is designed for the scenario where data journalists create serious content for the public. As a result, though SNIL can generate data articles for basketball games, it emphasizes the collaboration between AI-creators and data journalists as human-optimizers. Furthermore, in scenarios requiring creativity, the research of human-AI collaborative tools is also booming (e.g.(Chen et al., 2024; Xiao et al., 2023; Zhou et al., 2024a)). More research is needed to understand the relationship between the needs for different AI collaborators and the characteristics of scenarios. Then, in those scenarios requiring human-AI collaboration, designing suitable collaboration patterns should be studied. Finally, though these tools appears lack of human-AI collaboration from the data storytelling perspective, they support human-AI collaboration from a broader angle. For example, AI-created data stories help sports fans to understand tactics (Lee et al., 2024) and assist novices to comprehend charts (Ying et al., 2024). To conclude, we believe that human-AI collaboration remains an important topic to investigate and there will be even more opportunities to apply data storytelling for broader human-AI collaboration in data-related tasks.
5. Discussion
Section 4 presents exciting research progress in the era of generative AI, including new human-AI collaboration patterns and interactions in data storytelling tools. Meanwhile, we realize that it is not the end of this direction. There are plenty of needs and opportunities to continue our research, e.g., understanding these new patterns and investigating under-explored patterns (Figure 1(d)). This section provides suggestions for advancing this field with powerful AI models.
@@ 5. Discussion
Large-scale Generative AI Models as Universal Engines with Unified Interfaces. In addition to the benefits in Section 4.1, a more valuable characteristic of these generative AI models, especially LLMs, is to provide universal engines and interfaces for various roles and stages. Earlier tools often equip an AI collaborator with a specialized algorithm in one stage, e.g., fact mining and organization (Wang et al., 2019; Li et al., 2023b). Next, an interface, or “shared representation” (Heer, 2019) is required to support the communication among AI collaborators in different stages and even between AI and humans (Li et al., 2024b), e.g., chart or insight specification (Tang et al., 2017; Kim and Heer, 2021). Existing research shows that LLMs can support these stages as different roles with appropriate prompt design (Lee et al., 2024; Zhao et al., 2024). Furthermore, they can leverage natural languages as a unified interface for communication, with supplementary ways of programming languages. Therefore, these models can reduce tool development cost led by specialized algorithms and release the constraint of communication channels. Researchers can apply them to investigate expected but overlooked patterns, e.g., AI-optimizer + human-optimizer (Li et al., 2024b) and human-creator + AI-reviewer (Li et al., 2023a), conveniently. Furthermore, it will be exciting to explore fluid human-AI collaboration with user-configured patterns.
@@
Guidelines for Applying AI Models. Though powerful, these large-scale generative AI models have downsides, such as long response time (Li et al., 2025), and even risks like hallucination (Lee et al., 2023). As a result, their usage requires more research at a meta level to summarize when, where, and how they should be leveraged to achieve more benefit while mitigating their issues in data storytelling. For example, it is necessary to consider how to decompose tasks and design prompts (Shen et al., 2024a) and balance its usage with traditional algorithms to enhance the processing speed, accuracy, reliability, and etc. An design trade-off discussed by Socrate authors (Wu et al., 2023) is the usage of LLMs to extract intent for free-form feedback input by human-reviewers or only supporting limited input types but receiving more direct preferences. It will be interesting to propose AI application guidelines based on techniques, usage scenarios (Section 4.3), and collaboration patterns.
@@
Necessity for Continuous Reflection. In the generative AI era, we are witnessing a fast growth of data storytelling tools (Section 3.3). The research progress should be reflected continuously and more regularly to summarize lessons and identify opportunities. For example, Li et al. only discussed the existence of the relationship between collaboration patterns and usage scenarios with an example of AI-creators to support data story understanding (Li et al., 2024b). Powered by generative AI, more AI-creator-only scenarios emerges, allowing us to summarize their common characteristics, such as the casual and audience-oriented nature (Section 4.3). These findings greatly enrich our knowledge about the effect of usage scenarios on tool design. The example advises that human-AI collaboration evolves with technical advancements. Therefore, we must refresh our understanding continuously to gather useful experiences and keep them up-to-date.
6. Conclusion
The application of large-scale generative AI models has pushed forward the frontier of research in data storytelling tools. In this paper, we reflected the research progress through discussing the emerging and persistent human-AI collaboration patterns and their implications. We also identified the opportunities to advance this field. Due to space limit, this paper cannot discuss all exciting collaboration opportunities between humans and these AI models, such as applying video generation models (e.g., Sora (OpenAI, 2024)) for cinematic effects (Xu et al., 2023, 2022) or animated glyphs (Ying et al., 2022) and leveraging multi-agent systems (Shen et al., 2024a). Additionally, more roles of human and AI are worth exploration (Zhou et al., 2024b). We hope this paper can ignite future discussion about effective human-AI collaboration for data storytelling in the large-scale generative AI era.
Appendix A Visualizing the Trend of Data Storytelling Tools
Figure 2. The number of data storytelling tools between 2010 and 2024
[…]
This figure is a stacked bar chart to show the number of data storytelling tools between 2010 and 2024. The numbers across years are: 2 in 2010, 2 in 2013, 2 in 2014, 4 in 2016, 1 in 2017, 4 in 2018, 8 in 2019, 7 in 2020, 15 in 2021, 11 in 2022, 11 in 2023, and 20 in 2024. Among the 11 tools in 2023, 4 were collected by Li et al. and 7 were collected by us. The 20 tools in 2024 were fully collected by us.
Appendix B Visualizing the Relative Frequencies of Human-AI Collaboration Patterns
(a) Analysis
(b) Planning
(c) Implementation
(d) Communication
Figure 3. The human-AI collaboration pattern frequencies of earlier tools.
[…]
This figure shows the frequencies of collaboration patterns of earlier tools. The four subfigures show the pattern frequencies in four stages. In the analysis stage, the pattern of human-creator is 38%, human-creator and AI-assistant is 19%, AI-creator and human-assistant is 6%, AI-creator and human-optimizer is 25%, and AI-creator only is 12%. In the planning stage, the pattern of human-creator only is 52%, human-creator and AI-assistant is 10%, human-creator and AI-optimizer is 7%, AI-creator and human-optimizer is 28%, and AI-creator only is 3%. In the implementation stage, the pattern of human-creator only is 22%, human-creator and AI-assistant is 21%, AI-creator and human-assistant is 5%, AI-creator and human-optimizer is 34%, AI-assistant is 2%, AI-reviewer only is 3%, and AI-creator only is 12%. In the communication stage, the pattern of human-creator only is 60% and human-creator and AI-assistant is 40%. The other patterns that are not mentioned above have not been implemented in earlier tools.
(a) Analysis
(b) Planning
(c) Implementation
(d) Communication
Figure 4. The human-AI collaboration pattern frequencies of the latest tools.
[…]
This figure shows the frequencies of collaboration patterns of the latest tools. The four subfigures show the pattern frequencies in four stages. In the analysis stage, the pattern of human-creator is 14%, human-creator and AI-assistant is 21%, AI-creator and human-reviewer is 7%, AI-creator and human-optimizer is 21%, and AI-creator only is 36%. In the planning stage, the pattern of human-creator only is 35%, human-creator and AI-assistant is 12%, AI-creator and human-optimizer is 18%, AI-creator and human-reviewer is 6%, and AI-creator only is 29%. In the implementation stage, the pattern of human-creator only is 8%, human-creator and AI-assistant is 31%, AI-creator and human-optimizer is 27%, and AI-creator only is 35%. In the communication stage, the pattern of human-creator and AI-assistant is 33% and AI-creator only is 67%. The other patterns that are not mentioned above have not been implemented in the latest tools.

### v2

@@ 4.1 Which Patterns Are Still Widely Studied?
The most frequent patterns involving both humans and AI in earlier tools are human-creator + AI-assistant  [^4] The plus sign indicates the two roles collaborate in a specific stage. and AI-creator + human-optimizer since they allow humans to control the content while lowering the barrier of creation through automation [21]. In the latest tools, their frequency only has limited fluctuation, suggesting that they still attract considerable research interests (see Figs. 1 (a)-(c)). As the characteristics of the two patterns have been extensively discussed [21], we focus on the advancements brought by generative AI models when applying these patterns.
@@
First, large-scale generative AI models can improve the output and experience of human-AI collaboration. They can increase AI collaborators’ output space significantly. Several latest tools [46, 6, 41] apply AI-assistants to create visual components and provide design inspirations to support human-creators for authoring static or animated infographics. The similar design exists in earlier tools, such as DataQuilt [44]. However, earlier AI-assistants can only support re-using existing visual elements, greatly limiting the potential space of creation. The AI collaborators powered by retrieval-based algorithms also suffer from the issue, such as the AI-creator in Retrieve-then-Adapt [25]. Compared to earlier work, the latest tools powered by generative AI models are more flexible to create new visual elements beyond existing examples. Additionally, the interaction between collaborators can be more natural and closer to humans’ habits. In addition to prompting [41, 45], OutlineSpark [37] allows users to input a structured outline which guides AI-creator to generate slides. In Data Playwright [27], users write annotated narrations to specify expected animation effect and timing along with narrations. In both tools, users’ intent are conveniently expressed to enhance their agency, and their efforts are reduced through more AI automation. The last advantage is output quality improvement, such as enhancing text generation [24, 16].
Second, large-scale generative AI models have the potential to enable innovative application of AI collaborators. The previous usage of AI-assistant in the communication stage was to support individual users in live presentations, such as retrieving and presenting charts based on sketch [17]. Building on LLMs’ ability to understand and generate texts, Han and Isaacs [11] proposed an innovative usage case of AI-assistant to record the communication process between human-creators around charts and create an interactive data document based on the recording. The work demonstrates a new collaborative data story creation mode, where humans freely express their insights and AI documents and organizes humans’ ideas into a more structured data story. It is interesting to examine the benefit of this collaboration mode and explore other potentials. For example, what about asking AI to interview humans and write data stories or generating data videos like talk shows based on humans’ conversations around data? Furthermore, from the human-AI collaboration perspective, this AI-assistant has a notable difference with others. It serves as a third-party collaborator to document multiple humans’ communication, while not collaborating with individual humans and contributing to their creation or presentation. We should further investigate the potentials of human-AI collaboration to support multiple humans’ work in data storytelling [8]. The provocation also aligns with a finding from Fig. 1. In Fig. 1, many underexplored patterns are without any human- or AI-creators, such as AI-optimizer + human-optimizer. With future tools following these patterns, AI and humans will often collaborate to support third-party creators who do not use the tool.
4.2 What New Patterns Have Been Explored?
@@
From Figs. 1 (a), (b), and (d), we identified new collaboration patterns that were explored: human-reviewers in the analysis and the planning stages and AI-creators in the communication stage.
@@
The collaboration pattern of AI-creator + human-reviewer has been discussed as a potential strategy to enhance AI automation and human agency [21]. Compared to human-optimizers, human-reviewers can control the data story with feedback on AI-created content and thus save the effort to revise the content manually. Socrates [40] allows human-reviewer to provide feedback to the direction of data analysis (e.g., what attributes in datasets to be included in the data story) and the narrative structure (e.g., highlight a data fact using the strategy of contrasting it with another one). Specifically, the AI-creator, powered by a heuristic-based algorithm, will generate questions and ask for human-reviewer to rate, compare, or select options. Then, the AI-creator can leverage the results to create a data story that aligns with the intent of human-reviewer. Though they do not explicitly compare the efforts spent with their tool and another with human-optimizer, i.e., Calliope [33], their results show that the created stories with Socrates match the intent better, which also implies that humans might save efforts for further adjustment. The study results preliminarily verify the advantages of human-reviewer to achieve more AI automation and more human agency than human-optimizer, when collaborating with an AI-creator. In addition to the authors’ suggestion of exploring free-form feedback, future research can further study human-reviewers, such as understanding their pros and cons.
@@
Another new pattern is the application of AI-creators for communicating data stories automatically. According to an interview study, data workers often prefer AI not to perform the creator role in the communication since they may diminish the underlying value of human-human communication, such as building trust [20]. Therefore, we are particularly interested in the design of these AI-creators in the latest tools. Sportify [18] and AiCommentator [1] leverage AI-creators to communicate AI-identified data insights in sports games to humans. Ying et al. [43] and Shi et al. [32] support chart and dataset understanding with AI-creators who communicate insights to humans through AI-created animations and narrations. Comparing these cases with the interview study, we notice that these AI-creators communicates AI-created stories rather than presenting human-created stories on behalf of humans. It implies whether AI-creators can be accepted by humans in the communication stage may depend on AI’s roles in the previous stages. If the story is fully created by the AI collaborator in previous stages and reflect little human opinions, the role of AI-creators is more acceptable. On the other hand, when the story carries intensive information from human storytellers, the usage of AI-creators as a human proxy might raise concerns. Furthermore, the observation also indicates that the application of different roles across stages can be dependent on each other, which needs further investigation.
4.3 Which Patterns Are Gaining More Attention?
@@
In Fig. 1, we notice that the AI-creator-only pattern attracts growing research interests, where fully automatic tools across multiple stages emerge [43, 18, 28, 1]. We believe it is led by the increasing capability of large-scale generative AI models, such as the power of analyzing data [15] and generating text explanation [9]. One might question if the trend contradicts with the argument that human-AI collaboration in data storytelling would gain more interests [21]. Furthermore, does the advancement of AI capability mean that we should try to automate everything with AI in data storytelling and thus the research on human-AI collaboration is no longer needed?
@@
Inspired by the discussion where collaboration patterns can be affected by tool usage scenarios [21], we dived into the scenarios of these fully automatic tools to figure out why they are designed to be AI-creators-only. In the usage scenarios of Sportify [18] and AiCommentator [1], sports fans issue targeted queries about matches for desired information, such as tactics or player performances. Then, AI collaborators create tailored and casual data stories based on the queries and communicate them to humans automatically. Compared to other scenarios, such as presenting data stories to the public, this casual and low-stake one emphasizes the need of story customization and real-time creation while tolerating that the quality might not be perfect.
@@
To meet the demands of customization and high responsiveness, the creation and communication of data stories can hardly be intervened by human-creators or human-assistants, who are likely to take longer time and higher financial cost to finish tasks [20]. Traditional AI algorithms may also have limited flexibility for those highly customized data stories. The emergence of large-scale AI models helps fill the gap due to their capability to process free-form input that describes audience goals and to generate and communicate data stories accordingly (Sec. 4.1). They can act as AI-creators to automate the workflow. Since the scenarios are casual and low-stake, careful examination and quality improvement by human-optimizers or human-reviewers might not be needed.
@@
The scenario of Live Charts [43] and the work by Shi et al. [32] impose similar requirements to storytellers. To facilitate chart understanding, the authors of these two papers propose to transform static charts into data stories with animations and narrations. The generated data stories are intended to help users interpret their charts more quickly, while polished design is considered a lower priority. Given the high cost of humans further, a fully automatic pipeline is more suitable. Another similar audience-oriented scenario is to enhance text-only data story understanding by generating visuals [21], where automatic tools [14, 26] are designed.
@@
After examining these cases, we can attempt to answer the question about the necessity and potential future research direction of human-AI collaboration. First, led by the outstanding capability of large-scale generative AI models, it is a natural and irresistible trend that the application of AI-creators will be wider. To cope with the trend, we should actively experiment with these models [21] and learn how to apply them smartly, such as effective prompt design [28]. Furthermore, we should realize that the application of AI-creators is not only related to their performance but the scenario as well. The scenarios of fully automatic tools share common features, including low-stake scenarios that require real-time generation of personalized stories. Additionally, these scenarios typically do not involve human-human communication (Sec. 4.2). In other cases, fully automatic tools might not be suitable. For example, in contrast to the aforementioned tools [18, 1], SNIL [7] is designed for the scenario where data journalists create serious content for the public. As a result, though SNIL can generate data articles for basketball games, it emphasizes the collaboration between AI-creators and data journalists as human-optimizers. Furthermore, in scenarios requiring creativity, the research of human-AI collaborative tools is also booming (e.g., [6, 41, 46]). More research is needed to understand the relationship between the needs for different AI collaborators and the characteristics of scenarios. Then, in those scenarios requiring human-AI collaboration, designing suitable collaboration patterns should be further studied. Finally, though these tools appears lack of human-AI collaboration from the data storytelling perspective, they support human-AI collaboration from a broader angle. For example, AI-created data stories help sports fans to understand tactics [18] and assist novices to comprehend charts [43]. To conclude our response to the questions, we believe that human-AI collaboration remains an important topic to investigate and there will be more opportunities to apply data storytelling for broader human-AI collaboration in data-related tasks.
5 Future Opportunities
@@
Large-scale Generative AI Models as Universal Engines with Unified Interfaces. In addition to the benefits in Sec. 4.1, a more valuable characteristic of these generative AI models, especially LLMs, is to provide universal engines and interfaces for various roles and stages. Earlier tools often equip an AI collaborator with a specialized algorithm in one stage, e.g., fact mining and organization [23]. Next, an interface, or “shared representation” [13] is required to support the communication among AI collaborators in different stages and even between AI and humans [21], e.g., chart or insight specification. Existing research shows that LLMs can support these stages as different roles with appropriate prompt design [18]. Furthermore, they can leverage natural languages as a unified interface for communication, with supplementary ways of programming languages. Therefore, these models can reduce tool development cost led by specialized algorithms and release the constraint of communication channels. Researchers can apply them to investigate expected but overlooked patterns, e.g., AI-optimizer + human-optimizer [21] and human-creator + AI-reviewer [20], conveniently. Furthermore, it is exciting to explore fluid human-AI collaboration with user-configured patterns.
@@
Guidelines for Applying AI Models. Though powerful, these large-scale generative AI models have downsides, such as long response time and risks like hallucination. As a result, their usage requires more research at a meta level to summarize when, where, and how they should be leveraged to achieve more benefit while mitigating their issues. For example, it is necessary to consider how to decompose tasks and design prompts [28] and combine their usage with traditional AI algorithms to enhance the processing speed, accuracy, and reliability [22]. A design trade-off discussed by Socrates authors [40] is whether to apply LLMs for extracting intent from free-form feedback input by human-reviewers or only to support pre-defined input types for more direct preferences. It is interesting to propose AI usage guidelines based on techniques, usage scenarios, and collaboration patterns.
@@
Necessity for Continuous Reflection. In the generative AI era, we are witnessing a fast growth of data storytelling tools. The research progress should be reflected continuously and regularly to summarize lessons and identify opportunities. For example, Li et al. only discussed the existence of the relationship between collaboration patterns and usage scenarios with an example of AI-creators to support data story understanding [21]. Powered by generative AI, more AI-creator-only scenarios emerges, allowing us to summarize their common characteristics, such as the casual and audience-oriented nature (Sec. 4.3). These findings greatly enrich our knowledge about the effect of usage scenarios on tool design. Furthermore, the advancement of large-scale generative AI models greatly diversifies methods to leverage them. Users can finish various tasks in data storytelling with simple chat-based interfaces instead of carefully designed tools. For example, with commercial chatbots powered by multi-modal LLMs, such as GPT-4o and Gemini, users can analyze data, create texts, generate visuals, and iterate designs by prompting models. These observations advise that human-AI collaboration evolves with technical advancements. As a result, we must refresh our understanding continuously to gather useful knowledge related to both academic and commercial tools and keep it up-to-date.
Acknowledgements.
The authors thank Leixian Shen and all reviewers for suggestions.